### v1

@@ Abstract
Reconstructing full fields from extremely sparse and random measurements is a longstanding ill-posed inverse problem. A powerful framework for addressing such challenges is hierarchical probabilistic modeling, where uncertainty is represented by intermediate variables and resolved through marginalization during inference. Inspired by this principle, we propose Cascaded Sensing (Cas-Sensing), a hierarchical reconstruction framework that integrates an autoencoder–diffusion cascade. First, a neural operator-based functional autoencoder reconstructs the dominant structures of the original field—including large-scale components and geometric boundaries—from arbitrary sparse inputs, serving as an intermediate variable. Then, a conditional diffusion model, trained with a mask-cascade strategy, generates fine-scale details conditioned on these large-scale structures. To further enhance fidelity, measurement consistency is enforced via the manifold constrained gradient based on Bayesian posterior sampling during the generation process. This cascaded pipeline substantially alleviates ill-posedness, delivering accurate and robust reconstructions. Experiments on both simulation and real-world datasets demonstrate that Cas-Sensing generalizes well across varying sensor configurations and geometric boundaries, making it a promising tool for practical deployment in scientific and engineering applications.
@@
Field reconstruction; Probabilistic modeling; Cascade pipeline; Functional autoencoder; Diffusion model.
@@ 1 Introduction
Despite advances in sensor technology enabling large-scale deployments in physical systems, the number of monitoring sites is often constrained by high setup and maintenance costs. Therefore, spatial field reconstruction from limited local sensor information is a major challenge of complex physical systems, such as astrophysics [akiyama2019first], geophysics [kondrashov2006spatio, carrassi2018data, manohar2018data], atmospheric science [alonso2010novel, mishra2014compressed], and fluid dynamics [vinuesa2023transformative, buzzicotti2023data]. Due to the limited sensors, the reconstruction of multi-scale full-field data from extremely sparse and random observations is inherently an ill-posed problem, where the results often suffers from issues related to existence, uniqueness, and stability [hadamard2014lectures, breiding2023algebraic].
@@
Compressed sensing has been widely applied to address this challenge. It relies on the assumption that signals are sparse in a specific transform domain, and reconstructs full-field data from undersampled observations by solving optimization problems [candes2006robust]. To enhance reconstruction performance, compressed sensing variants combined with deep learning have been proposed, aiming to automatically learn effective sparse representations and reconstruction mappings in a data-driven manner [sun2017intelligent, shi2019image, ni2020deep, xu2022sparse]. While these methods have achieved certain theoretical and experimental progress, several limitations remain. Classic compressed sensing heavily depends on the sparsity prior, leading to performance degradation when dealing with non-sparse structures. Deep learning-based variants, although more flexible, are still often restricted to single-scale, fixed-domain reconstruction and struggle to capture multi-scale information.
@@
Besides the compressed sensing-based methods, many existing methods attempt to directly establish an end-to-end mapping between sparse measurements and the corresponding full-field data [fukami2021global, fan2025vitae, li2025novel, ghazijahani2025spatial]. However, such mappings are typically trained with a specific reorganization form of sparse data (for example, patch or voronoi tessellation) and geometric boundary. When encountering unseen sparse measurements—such as data acquired from a different sensor configuration—or when facing new geometric boundaries, these models often fail to produce accurate reconstructions and require retraining to adapt to the new scenarios. As a result, the applicability of these methods is generally limited to controlled or narrowly defined settings. They often struggle with generalization when applied to complex, real-world environments. Moreover, when the available observations become extremely sparse, the inherent ill-posed nature of the reconstruction problem severely impedes the ability to establish a stable and reliable mapping from sparse inputs to the full-field data [bell1978solutions, candes2011tight, arridge2019solving].
@@
Generative models, such as generative adversarial networks [goodfellow2020generative] and diffusion models [ho2020denoising, song2020score], have shown remarkable capabilities in producing high-quality, diverse data, with broad applications in vision [wang2023dr2, miao2024waveface, zhao2023towards], biomedicine [cao2024high, yu2020medical, nguyen2024sequence], and scientific computing [lienen2306zero, ruhling2023dyffusion, li2024synthetic]. By learning data distributions rather than explicit input–output mappings, these models can generate unseen but statistically consistent samples. Early studies applying generative models to field reconstruction primarily follow two approaches. One line of work learns supervised conditional distributions between full-field low-resolution and high-resolution data pairs [deng2019super, shu2023physics]. However, due to the immense variability of sparse measurement patterns, directly modeling the distribution of full-field data conditioned on sparse observations is practically infeasible, as it is impossible to cover all sparse-to-full field mappings during training. Another approach adopts a reverse diffusion step followed by a projection-based measurement consistency step [choi2021ilvr, chung2022come, kadkhodaie2021stochastic] or a Bayesian sampling step [du2024conditional, li2024learning] after pretraining an unconditional diffusion model to ensure the generated samples match sparse observations. However, this enforced measurement consistency often throws the sample path off the data manifold, leading to incorrect reconstructions [chung2022improving]. Although the manifold constrained gradient [chung2022improving] is proposed to mitigate this problem, under extremely sparse observations, the model still tends to produce random samples due to the overly loose constraints, resulting in non-unique and unreliable reconstructions. Similarly, FunDiff [wang2025fundiff] also relies on incorporating measurement consistency during the sampling stage to enforce conditional generation, and therefore suffers from degraded reconstruction quality under extremely sparse observations. Moreover, it extends diffusion models to function spaces by coupling a function autoencoder with latent diffusion, enabling physics-informed generative modeling across different discretizations. While powerful for capturing global structures, this latent-space generation strategy introduces an information bottleneck: fine-scale details may be lost during encoding and cannot be fully recovered during decoding, which can limit its effectiveness in detail reconstruction from sparse observations.
@@
Cascading pipeline of generative model is an effective way to improve the generation quality in case without strong conditioning information [nichol2021improved, saharia2022image, ho2022cascaded]. In computer vision, this approach involves training a sequence of models at progressively higher resolutions: a base model generates low-resolution samples, which are then refined by super-resolution models [ho2022cascaded]. Compared to training a single high-resolution model, cascading allocates most model capacity to low resolutions, which are critical for overall sample quality, while significantly improving training and inference efficiency. Moreover, individual models can be optimized independently, with architecture choices tailored to each resolution for optimal pipeline performance.
@@
Inspired by the cascading pipeline, the Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade is proposed to achieve accurate and robust multi-scale fields reconstruction from extremely sparse measurements. In Cas-Sensing, by viewing the full-field data as functional data, an neural operator-based functional autoencoder is first trained [bunker2024autoencoders] to extract the latent features of data, thus having the ability of robustly recovering the principal structures (large-scale components with geometric boundaries) from arbitrary sparse measurements. Then a conditional diffusion model is trained with the conditions—the reconstructed principal structures—to only focus on generating the small-scale details which are the difference between ground truth and conditions. To enhance the robustness of diffusion model, we introduce a Mask-Cascade Training (MCT) approach, which applies different random masks during training to generate diverse conditions from a fixed ground-truth field. After training the conditional diffusion model, we employ the Manifold Constrained Gradient (MCG) [chung2022improving] during sampling to ensure that the generated full-field data not only matches the sparse observations at measurement locations, but also remains consistent with the underlying data manifold. The proposed cascade pipeline effectively alleviates the ill-posedness of the original reconstruction problem by progressively recovering multi-scale components in a staged manner, thereby enabling Cas-Sensing to achieve high accuracy, strong generalizability, and enhanced robustness in reconstruction tasks.
@@ 2.1 Problem formulation
The problem of recovering the multi-scale full field $\boldsymbol{u}\in\mathbb{R}^{m}$ from a vector of sparse sensor measurements $\boldsymbol{y}\in\mathbb{R}^{n}$ can be formulated as modeling the conditional distribution $p(\boldsymbol{u}|\boldsymbol{y})$. However, when the observations are excessively sparse, directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$ is ill-posed, as the solution becomes highly non-unique. To address this challenge, we propose Cas-Sensing, a cascaded framework that mitigates the ill-posedness by leveraging the following marginalization rule:
@@
where $\boldsymbol{m}$ represents the principal structure (the large-scale component) of $\boldsymbol{u}$. This formulation naturally decomposes the task into two subtasks: modeling $p(\boldsymbol{m}|\boldsymbol{y})$ to infer the principal structure from sparse observations, and modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ to reconstrut the full field $\boldsymbol{u}$ by adding high-frequency details to $\boldsymbol{m}$ based on both $\boldsymbol{m}$ and $\boldsymbol{y}$. Since the principal structures of the signal are often sparse, reconstructing $\boldsymbol{m}$ from sparse observations—i.e., modeling $p(\boldsymbol{m}|\boldsymbol{y})$—is generally well-posed. Furthermore, compared to directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$, modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ incorporates an additional condition, constraining the reconstruction of $\boldsymbol{u}$ to remain consistent with the principal structures $\boldsymbol{m}$. Consequently, the degree of ill-posedness is substantially alleviated through this cascaded pipeline.
@@
In the proposed Cas-Sensing framework based on an autoencoder-diffusion cascade, a functional autoencoder is first trained to approximately model $p(\boldsymbol{m}|\boldsymbol{y})$. Then, a conditional diffusion model trained with the proposed mask-cascade training strategy to approximate $\mathbb{E}_{\boldsymbol{m}\sim p(\boldsymbol{m}|\boldsymbol{y})}[\,p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,]$. Specifically, once the autoencoder is trained and fixed, random sparse measurements $\boldsymbol{y}$ are generated from $\boldsymbol{u}$ and mapped through the autoencoder to produce corresponding $\boldsymbol{m}$. The conditional diffusion model is then trained to generate $\boldsymbol{u}$ conditioned on both $\boldsymbol{m}$ and $\boldsymbol{y}$, effectively learning to reconstruct the full field from varying sparse observations.
2.2 Functional autoencoder for principal structure reconstruction
@@
In practical engineering, sensors are often randomly placed, and many physical fields involve complex geometric boundaries, making field reconstruction from unstructured data challenging [chai2020deep]. To address this, we adopt a neural operator-based functional autoencoder [bunker2024autoencoders], which enables coordinate-based input for neural networks[regazzoni2024learning, krishnapriyan2023learning]. The functional autoencoder extracts latent features from unstructured data and robustly reconstruct the principal structure from random and sparse inputs.
@@
We model the inverse mapping from sparse observations $\boldsymbol{y}$ to the principal structure $\boldsymbol{m}$ using a deterministic autoencoder rather than a probabilistic model (e.g., a variational autoencoder). This choice is justified by the low uncertainty inherent in reconstructing $\boldsymbol{m}$ from $\boldsymbol{y}$ due to: (1) the strong sparse prior of $\boldsymbol{m}$ constraining the solution space since large-scale patterns vary smoothly and can be represented with few basis functions; (2) average pooling operation in the autoencoder promoting invariant and robust feature learning; and (3) our masked training strategy explicitly enhancing model’s robustness to noise and varying sensing patterns in $\boldsymbol{y}$. As a result, the posterior $p(\boldsymbol{m}|\boldsymbol{y})$ is highly concentrated, allowing us to approximate it as $p(\boldsymbol{m}|\boldsymbol{y})\approx\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$, where $\delta(\cdot)$ is a Dirac delta function and $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is the autoencoder output. This avoids instability from learning a full posterior variance, and still achieves accurate and robust reconstruction, as confirmed in our experiments.

### v2

@@ Abstract
Reconstructing full fields from extremely sparse and random measurements constitutes a fundamentally ill-posed inverse problem, in which deterministic end-to-end mappings often break down due to intrinsic non-uniqueness and uncertainty. Rather than treating sparse reconstruction as a regression task, we recast it as a hierarchical probabilistic inference problem, where uncertainty is explicitly represented, structured, and progressively resolved. From this perspective, we propose Cascaded Sensing (Cas-Sensing) as a general reconstruction paradigm for multi-scale physical fields under extreme data sparsity. Central to this paradigm is the introduction of an explicit intermediate representation that decomposes the original ill-posed problem into two substantially better-conditioned subproblems. First, a lightweight neural-operator-based functional autoencoder infers a coarse-scale approximation of the target field from sparse observations acting as an explicit intermediate variable. Rather than modeling multiple scales jointly, this intermediate estimate is deterministically fixed and subsequently used as the sole conditioning input to a conditional diffusion model that generates refined-scale details, yielding a cascaded inference structure with clearly separated reconstruction responsibilities. To ensure robustness under diverse sensing patterns, the diffusion model is trained using a mask-cascade strategy, which exposes it to a distribution of imperfect conditioning structures induced by extreme sparsity. During inference, measurement consistency is enforced through manifold-constrained gradients within a Bayesian posterior framework, ensuring fidelity to sparse observations while preserving data manifold coherence. This cascaded probabilistic formulation substantially alleviates ill-posedness, enabling accurate and stable reconstructions even under extreme sparsity. Extensive experiments on both simulated and real-world datasets demonstrate that Cas-Sensing generalizes effectively across sensor layouts and sparsity levels. These results suggest that Cas-Sensing provides not only an effective reconstruction tool, but also a principled and extensible framework for data-driven scientific sensing and inference in complex physical systems.
@@
Full-Field reconstruction; Probabilistic modeling; Cascade pipeline; Functional autoencoder; Diffusion model.
@@ 1 Introduction
Despite advances in sensor technology enabling coarse-scale deployments in physical systems, the number of monitoring sites is often constrained by the high cost of setup and maintenance. Therefore, spatial field reconstruction from limited local sensor information is a major challenge of complex physical systems, such as astrophysics [akiyama2019first], geophysics [kondrashov2006spatio, carrassi2018data, manohar2018data], atmospheric science [alonso2010novel, mishra2014compressed], and fluid dynamics [vinuesa2023transformative, buzzicotti2023data]. Due to the deployment of limited sensors, the reconstruction of multi-scale full-field data from extremely sparse and random observations is inherently an ill-posed problem with high uncertainty [hadamard2014lectures, breiding2023algebraic].
@@
Compressed sensing has been widely applied to address this type of problems. It relies on the assumption that signals are sparse in a specific transform domain such that uncertainty is suppressed, and reconstructs full-field data from undersampled observations by solving optimization problems [candes2006robust]. To enhance reconstruction performance, compressed sensing variants combined with deep learning have been proposed, aiming to automatically learn effective sparse representations and reconstruction mappings in a data-driven manner [sun2017intelligent, shi2019image, ni2020deep, xu2022sparse]. While these methods have achieved certain theoretical and experimental progress, several limitations remain. Classic compressed sensing heavily depends on the sparsity prior, leading to performance degradation when dealing with non-sparse structures. Deep learning-based variants, although more flexible, are still often restricted to single-scale, fixed-domain reconstruction and struggle to capture multi-scale information.
@@
Besides the compressed sensing-based methods, many existing methods attempt to directly establish an end-to-end mapping between sparse measurements and the corresponding full-field data [fukami2021global, fan2025vitae, li2025novel, ghazijahani2025spatial]. However, such mappings are typically trained with a specific reorganization form of sparse data (for example, patch or voronoi tessellation) and geometric boundary. When encountering unseen sparse measurements—such as data acquired from a different sensor deployment configuration, or when facing new geometric boundaries, these models often fail to produce accurate reconstructions and require retraining to adapt to the new scenarios. As a result, the applicability of these methods is generally limited to controlled or narrowly defined settings. They often struggle with generalization when applied to complex, real-world environments. From an inverse problem perspective, the limitation of end-to-end mappings is fundamentally rooted in the intrinsic uncertainty of sparse reconstruction. Under severely incomplete observations, the mapping from sparse measurements to full-field solutions is generally non-unique, meaning that multiple physically plausible fields may correspond to the same set of measurements. This ambiguity is a defining characteristic of ill-posed inverse problems and cannot be resolved by deterministic regression alone [bell1978solutions, arridge2019solving]. As a result, learning a single deterministic mapping inevitably collapses the underlying solution space, often leading to instability, overfitting to specific sensing configurations, and poor generalization when observations become sparser or deviate from the training distribution.
@@
In contrast, probabilistic generative modeling provides a natural mechanism for representing and sampling from the conditional distribution of feasible solutions, rather than committing to a single point estimate. By modeling the data generation process itself, generative models are able to capture uncertainty and variability arising from incomplete observations, making them particularly suitable for highly underdetermined reconstruction tasks. Generative models, such as generative adversarial networks [goodfellow2020generative] and diffusion models [ho2020denoising, song2020score], have shown remarkable capabilities in producing high-quality, diverse data, with broad applications in vision [wang2023dr2, miao2024waveface, zhao2023towards], biomedicine [cao2024high, yu2020medical, nguyen2024sequence], and scientific computing [lienen2306zero, ruhling2023dyffusion, li2024synthetic]. By learning data distributions rather than explicit input–output mappings, these models can generate unseen but statistically consistent samples. Early studies applying generative models to field reconstruction primarily follow two approaches. One line of work learns supervised conditional distributions between full-field low-resolution and high-resolution data pairs [deng2019super, shu2023physics]. However, due to the immense variability of sparse measurement patterns, directly modeling the distribution of full-field data conditioned on sparse observations is practically infeasible, as it is impossible to numerate all sparse-to-full field mappings during training. Another approach adopts a reverse diffusion step followed by a projection-based measurement consistency step [choi2021ilvr, chung2022come, kadkhodaie2021stochastic] or a Bayesian sampling step [du2024conditional, li2024learning] after pretraining an unconditional diffusion model to ensure the generated samples match sparse observations. However, this enforced measurement consistency often throws the sample path off the data manifold, leading to incorrect reconstructions [chung2022improving]. Although the manifold constrained gradient [chung2022improving] is proposed to mitigate this problem, under extremely sparse observations, the model still tends to produce random samples due to the overly loose constraints, resulting in non-unique and unreliable reconstructions. Similarly, FunDiff [wang2025fundiff] also relies on incorporating measurement consistency during the sampling stage to enforce conditional generation, and therefore suffers from degraded reconstruction quality under extremely sparse observations. Moreover, it extends diffusion models to function spaces by coupling a function autoencoder with latent diffusion, enabling physics-informed generative modeling across different discretizations. While this latent-space formulation is effective for capturing global structures, the unified generation of multi-scale components within a single latent representation constrains the model’s ability to independently regulate fine-scale details, particularly under extremely sparse conditioning.
@@
A cascade pipeline of generative models is an effective way to improve the generation quality in cases without strong conditioning information [nichol2021improved, saharia2022image, ho2022cascaded]. In computer vision, this approach involves training a sequence of models at progressively higher resolutions: a base model generates low-resolution samples, which are then refined by super-resolution models [ho2022cascaded]. Compared to training a single high-resolution model, cascading allocates most model capacity to low resolutions, which are critical for overall sample quality, while significantly improving training and inference efficiency. Moreover, individual models can be optimized independently, with architecture choices tailored to each resolution for optimal pipeline performance.
@@
As shown in Fig. 1, inspired by the hierarchical probabilistic decomposition via marginalization, the Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade is proposed to achieve accurate and robust multi-scale fields reconstruction from extremely sparse measurements. In Cas-Sensing, by viewing the full-field data as functional data (data sampled from continuous functions), a neural operator-based functional autoencoder is first trained [bunker2024autoencoders] to extract the latent features of data, thus having the ability of robustly recovering the coarse-scale structures from arbitrary sparse measurements. Then a conditional diffusion model is trained with the conditions—the reconstructed coarse-scale structures—to only focus on generating the fine-scale details, which are the residuals between ground truth and conditions. To enhance the robustness of diffusion model, we introduce a Mask-Cascade Training (MCT) approach, which applies different random masks during training to generate diverse conditions from a fixed ground-truth field. After training the conditional diffusion model, we employ the Manifold Constrained Gradient (MCG) [chung2022improving] during sampling to ensure that the generated full-field data not only matches the sparse observations at measurement locations, but also remains consistent with the underlying data manifold. The proposed cascade pipeline effectively alleviates the ill-posedness of the reconstruction problem by progressively recovering multi-scale components in a staged manner, thereby enabling Cas-Sensing to achieve high accuracy, strong generalizability, and enhanced robustness in reconstruction tasks.
@@ 2.1 Problem formulation
Figure 1: The overview of Cascaded Sensing, where $\boldsymbol{d}=\boldsymbol{u}-\boldsymbol{m}$ and MCG denotes Manifold Constrained Gradient for imposing the measurement consistency.
The problem of recovering the multi-scale full field $\boldsymbol{u}\in\mathbb{R}^{m}$ from a vector of sparse sensor measurements $\boldsymbol{y}\in\mathbb{R}^{n}$ can be formulated as modeling the conditional distribution $p(\boldsymbol{u}|\boldsymbol{y})$. However, when the observations are excessively sparse, directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$ is ill-posed, as the solution becomes highly non-unique. To address this challenge, we introduce an intermediate variable $\boldsymbol{m}$ into this probabilistic framework, thus forming a cascaded pipeline. Leverage the following hierarchical probabilistic decomposition via marginalization:
@@
where $\boldsymbol{m}$ represents the coarse-scale structure (the coarse-scale component) of $\boldsymbol{u}$. This formulation naturally decomposes the task into two subtasks: modeling $p(\boldsymbol{m}|\boldsymbol{y})$ to infer the coarse-scale structure $\boldsymbol{m}$ from sparse observations $\boldsymbol{y}$; and modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ to further reconstrut the full field $\boldsymbol{u}$ by adding refined details to $\boldsymbol{m}$, conditioned on both $\boldsymbol{m}$ and $\boldsymbol{y}$. At coarse scales, physical fields are dominated by low-frequency, smoothly varying components governed by global constraints, which can be represented with far fewer degrees of freedom than fine-scale fluctuations. As a result, reconstructing $\boldsymbol{m}$ from sparse observations –i.e., modeling $p(\boldsymbol{m}|\boldsymbol{y})$ – is generally well-posed. Furthermore, compared to directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$, modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ incorporates an additional condition, constraining the reconstruction of $\boldsymbol{u}$ to remain consistent with the coarse-scale structure $\boldsymbol{m}$. Consequently, the degree of ill-posedness is substantially alleviated through this cascaded pipeline.
@@
In the proposed Cas-Sensing framework based on an autoencoder-diffusion cascade, a functional autoencoder is first trained to approximately model $p(\boldsymbol{m}|\boldsymbol{y})$. Then, a conditional diffusion model trained with the proposed mask-cascade training strategy to approximate $\mathbb{E}_{\boldsymbol{m}\sim p(\boldsymbol{m}|\boldsymbol{y})}[\,p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,]$. Specifically, once the autoencoder is trained and fixed, random sparse measurements $\boldsymbol{y}$ are mapped through the autoencoder to produce corresponding $\boldsymbol{m}$. The conditional diffusion model is then trained to generate $\boldsymbol{u}$ conditioned on both $\boldsymbol{m}$ and $\boldsymbol{y}$, effectively learning to reconstruct the full field from varying sparse observations.
2.2 Functional autoencoder for coarse-scale structure reconstruction
Figure 2: The masked training scheme of functional autoencoder. The trained functional autoencoder is to approximate $p(\boldsymbol{m}|\boldsymbol{y})$ with $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$ as it learns the deterministic mapping between $\boldsymbol{y}$ and $\boldsymbol{m}$. $x_{e}$ and $u_{e}$ denote the coordinates and field values fed into the encoder, while $x_{d}$ and $u_{d}$ represent the coordinates and reconstructed values produced by the decoder.
@@ 2.2 Functional autoencoder for coarse-scale structure reconstruction
In practical engineering, sensors are often randomly placed (or some sensors are even movale), and many physical fields involve complex geometric boundaries, making field reconstruction from unstructured data challenging [chai2020deep]. To address this, we adopt a neural operator-based functional autoencoder [bunker2024autoencoders], which enables coordinate-based input for neural networks [regazzoni2024learning, krishnapriyan2023learning]. The functional autoencoder extracts latent features from unstructured data and robustly reconstruct the coarse-scale structure from random and sparse inputs.
@@
We model the inverse mapping from sparse observations $\boldsymbol{y}$ to the coarse-scale structure $\boldsymbol{m}$ using a deterministic autoencoder rather than a probabilistic model (e.g., a variational autoencoder). This choice is justified by the low uncertainty inherent in reconstructing $\boldsymbol{m}$ from $\boldsymbol{y}$ due to: (1) the strong sparse prior of $\boldsymbol{m}$ constraining the solution space since coarse-scale patterns vary smoothly and can be represented with few basis functions; (2) an average pooling operation in the autoencoder promoting invariant and robust feature learning; and (3) our masked training strategy explicitly enhancing model’s robustness to noise and varying sensing patterns in $\boldsymbol{y}$. As a result, the posterior $p(\boldsymbol{m}|\boldsymbol{y})$ is highly deterministic, allowing us to approximate it as:
$p(\boldsymbol{m}|\boldsymbol{y})\approx\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$ (2)
where $\delta(\cdot)$ is a Dirac delta function and $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is the autoencoder output. This avoids instability from learning a full posterior variance, and still achieves accurate and robust reconstruction, as confirmed in our experiments.